# On the complex zeros and the computational complexity of approximating the reliability polynomial

### v1

@@ 3.2. Relating the activity-locus and the the density locus
Let $G_{0}$ be a two-terminal graph and suppose that $p\in\mathcal{A}_{G_{0}}\cup\mathcal{A}^{\mathbb{R}}_{G_{0}}$. Let $g$ be the Möbius transformation as defined in (3.2). Then $g(1)=1$ and $g(p)=p$ and $1$ is an attracting fixed point of $g$, while $p$ is a repelling fixed point of $g$.
@@
If $p\in\mathbb{R}$ we have $\hat{y}_{G_{0}}(p)<-1$ and hence $y_{G_{n}}-1$ alternates in sign. Therefore there exists $n\geq m_{\varepsilon}$ such that $R(G_{n};p)R(G_{n+1};p)\neq 0$, $\hat{y}_{G_{0}}(p)-1\in(-\varepsilon,0)$ and $\hat{y}_{G_{0}}(p)-1\in(0,\varepsilon)$. This implies that $\mathbb{N}(y_{G_{n}}-1)+\mathbb{N}(y_{G_{n+1}}-1)$ is $\varepsilon$-dense in $\mathbb{R}$ and by Lemma 2.3 consists of shifted effective edge interactions of series compositions of $G_{n}$ and $G_{n+1}$. By Lemma 2.1 $R(G_{n}^{\bowtie k}\bowtie G_{n+1}^{\bowtie\ell};p)\neq 0$ for all $k,\ell$. It follows that $\{y_{G}\mid G\in\mathcal{H}_{G_{0}},R(G;p)\neq 0\}$ is $\varepsilon$-dense in $\mathbb{R}$. As this holds for all $\varepsilon>0$, it follows that $p\in\mathcal{D}^{\mathbb{R}}_{G_{0}}$ in case $p$ is real, thus proving the inclusion $\mathcal{A}^{\mathbb{R}}_{G_{0}}\setminus\{p\mid R(G_{0};p)=0\}\subseteq\mathcal{D}^{\mathbb{R}}_{G_{0}}$.
@@ 4.1. Boundedness of reliability zeros versus activity near the positive real axis
Let $p\in U$. If $p\in\mathbb{R}$, then $1/\hat{y}_{G}(p)\in\mathbb{R}$ and hence is by definition not contained in $\mathcal{A}.$ Next suppose $p\notin\mathbb{R}$ and assume towards contradiction that for some two-terminal graphs $G,H$ we have $1/\hat{y}_{G}(p)\in\mathcal{A}_{H}$ and $p\notin\mathcal{E}(G)$. This means that $1<|\hat{y}_{H}(1/\hat{y}_{G}(p))|<\infty$ and moreover that $\hat{y}_{H}(1/\hat{y}_{G}(p))$ is not real. By Lemma 2.7 we know that $\hat{y}_{H}(1/\hat{y}_{G}(p))=\hat{y}_{G(H)}(p)$ and therefore $p$ is contained in $\mathcal{A}_{G(H)}$, since $p\notin\mathcal{E}_{G(H)}$ because (2.3) implies $R(G(H);p)+S(G(H);p)\neq 0$. This contradicts the fact that $U\cap\mathcal{A}=\emptyset$. ∎
@@
Let us prove (1) $\Rightarrow$ (2). Thus let us assume that $\mathcal{A}$ is not dense in $\mathbb{C}$, in particular there exists an open set $U\subseteq\mathbb{C}$, such that $U\cap\mathcal{A}=\emptyset$. As observed in the introduction we have that $\mathbb{D}\setminus\{0\}\subseteq\mathcal{A}$ (because for the two terminal graph $K_{2}$ we have $\hat{y}_{K_{2}}(p)=1/p)$. We may therefore assume that for any $p\in U$ we have $|p|>1$. Since $U$ is not empty and open, we can find an arc $\Gamma$ of a circle of radius $r>0$ with central angle $\theta>0$ in $U$ and such that there is a $\varepsilon>0$ such that the $\varepsilon$-neighborhood of $\Gamma$ is still part of $U$.
@@
By the previous lemma each $U_{m}$ is disjoint from $\mathcal{A}$, since the exceptional sets of the parallel compositions $K_{2}^{\parallel m}$ are empty. We claim that $U=\bigcup_{m\in\mathbb{N}}U^{\prime}_{m}\cup\{\infty\}$ will witness the boundedness of $\mathcal{A}$. First observe that if $m$ is sufficiently large, then $(r+\varepsilon)^{m}>(r-\varepsilon)^{m+1}$ and thus the union of the annuli $U^{\prime}_{m},U^{\prime}_{m+1}$ form a connected annulus. As $r+\varepsilon>1$, we obtain that $\cup_{m\in\mathbb{N}}U^{\prime}_{m}\supseteq\mathbb{C}\setminus B(0,M)$ for some $M>0$ and is disjoint from $\mathcal{A}$. This shows that $\mathcal{A}$ is a bounded set, proving (1) $\Rightarrow$ (2).
@@
The exceptional set of $K_{2}^{\bowtie_{n}}$ consists of those $p$ for which $R(K_{2}^{\bowtie_{n}};p)+S(K_{2}^{\bowtie_{n}};p)=0$, or equivalently those $p$ for which $S(K_{2}^{\bowtie_{n}};p)\neq 0$ and $\hat{y}_{K_{2}^{\bowtie_{n}}}(p)=0$ or $R(K_{2}^{\bowtie_{n}};p)=0=S(K_{2}^{\bowtie_{n}};P)=0$. Now since $R(K_{2}^{\bowtie_{n}};p)=R(K_{2};p)^{n}=(1-p)^{n}$ only has $p=1$ as zero and $\hat{y}_{K_{2}^{\bowtie_{n}}}(p)=\frac{1+(n-1)p}{np}$, by Lemma 2.4, it follows that
@@
For $n>M+2$ we have $\frac{nM}{1+nM-M}<\frac{n+1}{n}$ and therefore $U_{n}\cap\mathbb{R}$ and $U_{n+1}\cap\mathbb{R}$ have a nonempty intersection. Since $\frac{nM}{1+nM-M}\to 1$ as $n\to\infty$, it follows that $\bigcup_{n\in\mathbb{N}}U_{n}$ contains an interval of the form $(1,w)$ for some $w>1$. By (4.2) and (4.1) it follows that $\bigcup_{n\in\mathbb{N}}U_{n}$ contains an open set $O_{1}$ disjoint from each of the exceptional sets $\mathcal{E}(K_{2}^{\bowtie_{n}})$ such that $O_{1}$ contains $(1,w)$ and is disjoint from $\mathcal{A}$.
@@ 4.2. Activity locus and the unit circle
In this subsection we aim to show that all but finitely many points of the closed unit circle are part of the activity locus. To do so, let us recall the multivariate version of the edge replacement construction see Lemma 2.9. Let a template graph $(H,c)$ be a two-terminal graph $H$ with edge label $c:E(H)\to\{a_{1},a_{2}\}$. Then for any $G_{1},G_{2}$ two-terminal graph let $H(G_{1},G_{2})$ be the graph obtained by replacing each edge labeled by $a_{i}$ with $G_{i}$. This is a special case of the construction described in Lemma 2.9.
@@
If $G_{1},G_{2}$ are two-terminal graphs and $(H,c)$ is the pentagonal template from Figure 3, then
@@
To prove Proposition 3.2 we need the following technical lemma about $F(y_{1},y_{2})$ defined the previous lemma.
@@
Now let us assume that $\alpha$ is irrational. In what follows we use the fact that for any irrational angle $\beta$ the orbit $\{e^{k(2\pi i\beta)}\mid k\in\mathbb{N}\}$ is dense in the circle. Therefore there exists a $k\in\mathbb{N}$ such that $0\neq\arg(p^{-k})\in I$. Thus by the previous lemma $|F(p^{-k},p^{k})|>1$. Since $F(y_{1},y_{2})$ is a continuous function, there exists a $\delta>0$ such that $|F(p^{-k},y_{2})|>1$ for all $y_{2}\in B_{\delta}(p^{k})$. Since $k\alpha$ is irrational there exists $\ell\in\mathbb{N}$ such that $(p^{-k})^{\ell}\in B_{\delta}(p^{k})$. This means that $|f(p^{-k},p^{-k\ell})|>1$. Now let $G_{1}=K_{2}^{\parallel k}$ and $G_{2}=K_{2}^{\parallel k\ell}$, then
@@ 5. Density implies hardness
for an edge $e$ of $F$ and combine these in a telescoping fashion to compute $R(F;p)$ exactly. We do this by viewing the ratio as the solution $x^{*}$ to a linear equation of the form AxB=0Ax-B=0 with $B=R(F;p)$ and $A=R(F\setminus e;p)$ and $x=y_{G^{\prime}}-(p+1)$ for some $G^{\prime}\in\mathcal{H}_{G}$ such that $R(G^{\prime};p)\neq 0$. Making use of Lemma 2.2, we can view $Ax+B$ as the reliability polynomial of $F(G^{\prime})_{e}$ (the graph obtained by implementing $G^{\prime}$ on the edge $e$ of $F$) and use the oracle to approximately determine the value of $Ax+B$. By using different values of $x$ that we can achieve using the fact that $p$ is contained in the desnity locus of, $\in\mathcal{D}_{G}$, in combination with a form of binary search we can then determine $x^{*}$ with very high precision. This requires us to actually generate any given value $x$ with very high precision as $y_{G^{\prime}}-(p+1)$ for some $G^{\prime}\in\mathcal{H}_{G}$ such that $R(G^{\prime};p)\neq 0$ in polynomial time. Using the fact that algebraic numbers of bounded complexity form a discrete set (much like the rational numbers of bounded denominator) we can then determines the value $x^{*}$ exactly using an algorithm due to Kannan, Lenstra and Lovász [kannan-1988].
@@ 5.1. Representing algebraic numbers
We need a few more definitions. Let $q=\sum_{i=0}^{d}a_{i}x^{i}$ be a polynomial of degree $d$, leading coefficient $a_{d}$ and roots $\{\alpha_{i}\}_{i\in\{1,\ldots,d\}}$.
@@ 5.2. Exact Computation
$\text{Rel}(p)=\left(\frac{p-1}{p}\right)^{1-|V|}p^{|E|}T(G;1,1/p).$
Since connectedness can be verified in polynomial time by a simple search algorithm (Breadth First Search and Depth First Search both work), Planar-Rel(p)(p) reduces trivially to Exact-Tutte(1,1/p)(1,1/p) where for algebraic numbers $x,y$, Planar-Tutte(x,y)(x,y) is defined as follows.
@@ 5.3. Exponential density
Let $G_{0}$ be a two-terminal graph and let $p\in\mathcal{D}_{G_{0}}$ (resp. $p\in\mathcal{D}^{\mathbb{R}}_{G_{0}}$) such that $R(G_{0};p)\neq 0$ and $S(G_{0};p)\neq 0$. Then there exists an algorithm that on input of $y_{0}\in\mathbb{Q}[i]$ (resp. $y_{0}\in\mathbb{Q}$) and rational $\varepsilon>0$ outputs a two-terminal graph $G\in\mathcal{H}_{G_{0}}$ and the value $R(G;p)$ satisfying $|y_{G}(p)-(p+1)-y_{0}|<\varepsilon$ and $R(G;p)\neq 0$. Both the running time of the algorithm and the size of $G$ are ${\mathtt{poly}}({\mathtt{size}}(y_{0},\varepsilon))$.
@@
We call a Möbius transformation $\Phi$ is contracting on a set $U\subseteq\mathbb{C}$ if $|\Phi^{\prime}(z)|<1$ for all $z\in U$. We restate Lemma 2.8 from [Bezakovahardcore].
@@
Suppose we have $m\in\mathbb{Q}[i]$ and $r>0$ rational; let $U=B(m,r)$. Further suppose that we have Möbius transformations with algebraic coefficients $\Phi_{i}:\hat{\mathbb{C}}\to\hat{\mathbb{C}}$ for $i\in[\ell]$ satisfying the following:
@@
    $U\subseteq\bigcup_{i\in[t]}\varphi_{i}(U)$.
Then there is an algorithm which on input of algebraic numbers $s,t\in U$ (respectively the starting point and target) and rational $\varepsilon>0$ outputs in ${\mathtt{poly}}({\mathtt{size}}(s,t,\varepsilon))$-time an algebraic number $x^{\prime}\in B(x,\varepsilon)$ and a sequence $i_{1},\ldots,i_{k}\in[\ell]$ such that
@@
    $x=\Phi_{i_{k}}\circ\ldots\circ\Phi_{i_{1}}(s)$ and
@@
    $\Phi_{i_{j}}\circ\ldots\circ\Phi_{i_{1}}(s)\in U$ for all $j\leq k$.
@@
As remarked in [GalanisetalboundeddegreeIsing] the proof of [Bezakovahardcore]*Lemma 2.8 gives this directly. It moreover also holds in the special case when all parameters are real. ∎
@@
By Theorem 3.5 we have $p\in\mathcal{A}_{G_{0}}$ (resp. $p\in\mathcal{A}^{\mathbb{R}}_{G_{0}}$) and thus by Lemma 3.4 we have $|g^{\prime}(1)|<1$.
@@
Consider the open set $U_{1}=\{z\in\mathbb{C}\mid|g^{\prime}(z)|<1\}$. Clearly, $1\in U_{1}$. Since $U_{1}$ is open there exists $r>0$ such that $U:=B(1,r)\subseteq U_{1}$. (We can easily determine such $r$ explicitly from the formula for $g^{\prime}(z)$.) Then for every $u\in U$ we have that $z\mapsto g(z)+u-1$ is a contraction on $U$. Since $\{y_{G}(p)\mid G\in\mathcal{H}_{G_{0}},R(G;p)\neq 0\}$ is dense in $\mathbb{C}$ the collection of open sets of the form $g(U)+y_{G}(p)-1$ with $G\in\mathcal{H}^{*}_{G_{0}}(p)$ covers the compact set $\overline{U}$. Therefore there exists a finite set of graphs $G_{i}\in\mathcal{H}^{*}_{G_{0}}(p)$, $i=1,\ldots,t$ such the associated sets cover $\overline{U}$.
The corresponding set of contracting Möbius transformations $\{\Phi_{i}(z)=g(z)+y_{G_{i}}(p)-1\}_{i\in[t]}$ thus satisfies the hypothesis of Lemma 5.3. ∎
@@
Now we are ready to describe the desired algorithm. Let $\omega_{0}=y_{0}+(1+p)$. Our goal to find an effective edge interaction $\omega_{1}=y_{G}(p)$ such that $|\omega_{1}-\omega_{0}|\leq\varepsilon$ in polynomial time of the input. For the algorithm we consider three different cases: (1) $\omega_{0}\in U$, (2) $\omega_{0}\in B(p,2)\setminus U$ and (3) $\omega_{0}\in\mathbb{C}\setminus(U\cup B(p,2))$.

### v2

@@ 3.2. Relating the activity-locus and the the density locus
Let $G_{0}$ be a two-terminal graph and suppose that $p\in\mathcal{A}_{G_{0}}\cup\mathcal{A}^{\mathbb{R}}_{G_{0}}$. Let $g$ be the Möbius transformation as defined in (3.2). Then $g(1)=1$ and $g(p)=p$. Moreover, $1$ is an attracting fixed point of $g$, while $p$ is a repelling fixed point of $g$.
@@
If $p\in\mathbb{R}$ we have $\hat{y}_{G_{0}}(p)<-1$ and hence $y_{G_{n}}-1$ alternates in sign, because $g^{\prime}(1)=1/\hat{y}_{G_{0}}(p)<0$. Therefore there exists $n\geq m_{\varepsilon}$ such that $R(G_{n};p)R(G_{n+1};p)\neq 0$, $\hat{y}_{G_{n}}(p)-1\in(-\varepsilon,0)$ and $\hat{y}_{G_{n}+1}(p)-1\in(0,\varepsilon)$. This implies that $\mathbb{N}(y_{G_{n}}-1)+\mathbb{N}(y_{G_{n+1}}-1)$ is $\varepsilon$-dense in $\mathbb{R}$ and by Lemma 2.3 consists of shifted effective edge interactions of series compositions of $G_{n}$ and $G_{n+1}$. By Lemma 2.1 $R(G_{n}^{\bowtie k}\bowtie G_{n+1}^{\bowtie\ell};p)\neq 0$ for all $k,\ell$. It follows that $\{y_{G}\mid G\in\mathcal{H}_{G_{0}},R(G;p)\neq 0\}$ is $\varepsilon$-dense in $\mathbb{R}$. As this holds for all $\varepsilon>0$, it follows that $p\in\mathcal{D}^{\mathbb{R}}_{G_{0}}$ in case $p$ is real, thus proving the inclusion $\mathcal{A}^{\mathbb{R}}_{G_{0}}\setminus\{p\mid R(G_{0};p)=0\}\subseteq\mathcal{D}^{\mathbb{R}}_{G_{0}}$.
@@ 4.1. Boundedness of reliability zeros versus activity near the positive real axis
Let $p\in U$. If $p\in\mathbb{R}$, then $1/\hat{y}_{G}(p)\in\mathbb{R}$ and hence is by definition not contained in $\mathcal{A}.$ Next suppose $p\notin\mathbb{R}$ and assume towards contradiction that for some two-terminal graphs $G,H$ we have $1/\hat{y}_{G}(p)\in\mathcal{A}_{H}$ and $p\notin\mathcal{E}(G)$. This means that $1<|\hat{y}_{H}(1/\hat{y}_{G}(p))|<\infty$ and moreover that $\hat{y}_{H}(1/\hat{y}_{G}(p))$ is not real. By Lemma 2.7 we know that $\hat{y}_{H}(1/\hat{y}_{G}(p))=\hat{y}_{G(H)}(p)$ and therefore $p$ is contained in $\mathcal{A}_{G(H)}$, since $p\notin\mathcal{E}_{G(H)}$ because (2.3) implies $R(G(H);p)+{\mathsf{SP}}(G(H);p)\neq 0$. This contradicts the fact that $U\cap\mathcal{A}=\emptyset$. ∎
@@
Let us prove (1) $\Rightarrow$ (2). Thus let us assume that $\mathcal{A}$ is not dense in $\mathbb{C}$, in particular there exists an open set $U\subseteq\mathbb{C}$, such that $U\cap\mathcal{A}=\emptyset$. As observed in the introduction we have that $\mathbb{D}\setminus\{0\}\subseteq\overline{\mathcal{A}}$ (because for the two terminal graph $K_{2}$ we have $\hat{y}_{K_{2}}(p)=1/p)$. We may therefore assume that for any $p\in U$ we have $|p|>1$. Since $U$ is not empty and open, we can find an arc $\Gamma$ of a circle of radius $r>1$ with central angle $\theta>0$ in $U$ and such that there is $\varepsilon>0$ such that the $\varepsilon$-neighborhood of $\Gamma$ is still part of $U$.
@@
By the previous lemma each $U_{m}$ is disjoint from $\mathcal{A}$, since the exceptional sets of the parallel compositions $K_{2}^{\parallel m}$ are empty. We claim that $U^{\prime}=\bigcup_{m\in\mathbb{N}}U^{\prime}_{m}\cup\{\infty\}$ will witness the boundedness of $\mathcal{A}$. First observe that if $m$ is sufficiently large, then $(r+\varepsilon)^{m}>(r-\varepsilon)^{m+1}$ and thus the union of the annuli $U^{\prime}_{m},U^{\prime}_{m+1}$ form a connected annulus. As $r+\varepsilon>1$, we obtain that $\cup_{m\in\mathbb{N}}U^{\prime}_{m}\supseteq\mathbb{C}\setminus B(0,M)$ for some $M>0$ and is disjoint from $\mathcal{A}$. This shows that $\mathcal{A}$ is a bounded set, proving (1) $\Rightarrow$ (2).
@@
The exceptional set of $K_{2}^{\bowtie_{n}}$ consists of those $p$ for which $R(K_{2}^{\bowtie_{n}};p)+{\mathsf{SP}}(K_{2}^{\bowtie_{n}};p)=0$, or equivalently those $p$ for which ${\mathsf{SP}}(K_{2}^{\bowtie_{n}};p)\neq 0$ and $\hat{y}_{K_{2}^{\bowtie_{n}}}(p)=0$ or $R(K_{2}^{\bowtie_{n}};p)=0={\mathsf{SP}}(K_{2}^{\bowtie_{n}};P)=0$. Now since $R(K_{2}^{\bowtie_{n}};p)=R(K_{2};p)^{n}=(1-p)^{n}$ only has $p=1$ as zero and $\hat{y}_{K_{2}^{\bowtie_{n}}}(p)=\frac{1+(n-1)p}{np}$, by Lemma 2.4, it follows that
@@
This implies that the exceptional sets $\mathcal{E}(K_{2}^{\bowtie_{n}})$ are disjoint from $U$ and hence by (4.1) we have
(4.3) $\text{ for all $n$}\quad f_{n}(U)\cap\mathcal{A}=\emptyset.$
@@
For $n>M+2$ we have $\frac{nM}{1+nM-M}<\frac{n+1}{n}$ and therefore $U_{n}\cap\mathbb{R}$ and $U_{n+1}\cap\mathbb{R}$ have a nonempty intersection. Since $\frac{nM}{1+nM-M}\to 1$ as $n\to\infty$, it follows that $O_{1}:=\bigcup_{n\in\mathbb{N}}U_{n}$ contains an interval of the form $(1,w)$ for some $w>1$. By (4.3) $O_{1}$ is disjoint from $\mathcal{A}$.
@@ 4.2. Activity locus and the unit circle
In this subsection we aim to show that all but finitely many points of the closed unit circle are part of the activity locus. To do so, let us recall the multivariate version of the edge replacement construction see Lemma 2.9. Define a template graph $(H,c)$ to be a two-terminal graph $H$ with edge label $c:E(H)\to\{a_{1},a_{2}\}$. For any $G_{1},G_{2}$ two-terminal graphs let $H(G_{1},G_{2})$ be the graph obtained by replacing each edge labeled by $a_{i}$ with $G_{i}$. This is a special case of the construction described in Lemma 2.9.
@@
Let $G_{1},G_{2}$ be two two-terminal graphs and let $(H,c)$ be the pentagon template from Figure 3. Then
@@
To prove Proposition 3.2 we need the following technical lemma about $F(y_{1},y_{2})$ defined in the previous lemma.
@@
Now let us assume that $\alpha$ is irrational. In what follows we use the fact that for any irrational angle $\beta$ the orbit $\{e^{k(2\pi i\beta)}\mid k\in\mathbb{N}\}$ is dense in the unit circle. Therefore there exists a $k\in\mathbb{N}$ such that $0\neq\arg(p^{-k})\in I$. Thus by the previous lemma $|F(p^{-k},p^{k})|>1$. Since $F(y_{1},y_{2})$ is a continuous function, there exists a $\delta>0$ such that $|F(p^{-k},y_{2})|>1$ for all $y_{2}\in B_{\delta}(p^{k})$. Since $k\alpha$ is irrational there exists $\ell\in\mathbb{N}$ such that $(p^{-k})^{\ell}\in B_{\delta}(p^{k})$. This means that $|f(p^{-k},p^{-k\ell})|>1$. Now let $G_{1}=K_{2}^{\parallel k}$ and $G_{2}=K_{2}^{\parallel k\ell}$, then
@@ 5. Density implies hardness
for an edge $e$ of $F$ and combine these in a telescoping fashion to compute $R(F;p)$ exactly. We do this by viewing the ratio as the solution $x^{*}$ to a linear equation of the form AxB=0Ax-B=0 with $B=R(F;p)$ and $A=R(F\setminus e;p)$ and $x=y_{G^{\prime}}-(p+1)$ for some $G^{\prime}\in\mathcal{H}_{G}$ such that $R(G^{\prime};p)\neq 0$. Making use of Lemma 2.2, we can view $Ax+B$ as the reliability polynomial of $F(G^{\prime})_{e}$ (the graph obtained by implementing $G^{\prime}$ on the edge $e$ of $F$) and use the oracle to approximately determine the value of $Ax+B$. By using different values of $x$ that we can achieve using the fact that $p$ is contained in the density locus of $\mathcal{H}_{G}$, in combination with a form of binary search we can then determine $x^{*}$ with very high precision. This requires us to actually generate any given value $x$ with very high precision as $y_{G^{\prime}}-(p+1)$ for some $G^{\prime}\in\mathcal{H}_{G}$ such that $R(G^{\prime};p)\neq 0$ in polynomial time. Using the fact that algebraic numbers of bounded complexity form a discrete set (much like the rational numbers of bounded denominator) we can then determines the value $x^{*}$ exactly using an algorithm due to Kannan, Lenstra and Lovász [kannan-1988].
@@ 5.1. Representing algebraic numbers
We need a few more definitions. Let $q=\sum_{i=0}^{d}a_{i}x^{i}$ be a polynomial of degree $d$, with leading coefficient $a_{d}$ and roots $\alpha_{1},\ldots,\alpha_{d}$.
@@ 5.2. Exact Computation
$\text{Rel}(p)=\left(\frac{1-p}{p}\right)^{|V|-1}p^{|E|}T(G;1,1/p).$
Since connectedness can be verified in polynomial time by Breadth First Search for example, Planar-Rel(p)(p) reduces trivially to Planar-Tutte(1,1/p)(1,1/p), where for algebraic numbers $x,y$, Planar-Tutte(x,y)(x,y) is defined as follows.
@@ 5.3. Exponential density
Let $G_{0}$ be a two-terminal graph and let $p\in\mathcal{D}_{G_{0}}$ (resp. $p\in\mathcal{D}^{\mathbb{R}}_{G_{0}}$) such that $R(G_{0};p)\neq 0$ and $S(G_{0};p)\neq 0$. Then there exists an algorithm that on input of $y_{0}\in\mathbb{Q}[i]$ (resp. $y_{0}\in\mathbb{Q}$) and rational $\varepsilon>0$ outputs a two-terminal graph $G\in\mathcal{H}_{G_{0}}$ and the value $R(G;p)$ satisfying $|y_{G}(p)-(p+1)-y_{0}|<\varepsilon$ and $R(G;p)\neq 0$. Both the running time of the algorithm and the size of $G$ are ${\mathsf{poly}}({\mathsf{size}}(y_{0},\varepsilon))$.
@@
We call a Möbius transformation $\Phi$ contracting on a set $U\subseteq\mathbb{C}$ if there exists $\alpha\in(0,1)$ such that $|\Phi^{\prime}(z)|<\alpha$ for all $z\in U$. We restate a slightly modified version of Lemma 2.8 from [Bezakovahardcore].
@@
Let $m\in\mathbb{Q}[i]$, $r>0$ rational and $U=B(m,r)$. Suppose that we have Möbius transformations with algebraic coefficients $\Phi_{i}:\hat{\mathbb{C}}\to\hat{\mathbb{C}}$ for $i\in[\ell]$ satisfying the following:
@@
    $U\subseteq\bigcup_{i\in[\ell]}\Phi_{i}(U)$.
Then there is an algorithm which on input of algebraic numbers $s,t\in U$ (respectively the starting point and target) and rational $\varepsilon>0$ outputs in ${\mathsf{poly}}({\mathsf{size}}(s,t,\varepsilon))$-time an algebraic number $x\in B(t,\varepsilon)$ and a sequence $i_{1},\ldots,i_{k}\in[\ell]$ such that
@@
    $x=\Phi_{i_{1}}\circ\cdots\circ\Phi_{i_{k}}(s)$ and
@@
    $\Phi_{i_{j}}\circ\cdots\circ\Phi_{i_{k}}(s)\in B(m,3r)$ for all $j\leq k$.
@@
Remark 5.4.
We note that item (iii) in the lemma above is not part of the statement of [Bezakovahardcore]*Lemma 2.8. Additionally, we require the $\Phi_{i}$ to be contracting on $B(m,3r)$ rather than on $U=B(m,r)$. The reason for this restatement is that we believe that the proof given in [Bezakovahardcore] is not correct, since it implicitly assumes that all values $\Phi_{i_{j}}\circ\cdots\circ\Phi_{i_{k}}(s)$ are contained in $U$ to be able to apply the bound on the derivative. This however does not have to be the case. We note that our restatement can be applied in essentially the same way as it was originally applied in [Bezakovahardcore] and later in [GalanisetalboundeddegreeIsing] and [main-approx]. Therefore the error in the proof does not have any implications for the contents of these papers.
@@
We follow the ideas of the proof of [Bezakovahardcore]*Lemma 2.8 and modify some of the steps appropriately. Let $M\in(0,1)$ be an upper bound on $|\Phi^{\prime}(z)|$ for all $i=1,\ldots,\ell$ and $z\in B(m,3r)$. Set $t_{0}=t$, then recursively define $t_{j}$ for $j\geq 1$ by choosing an index $i_{j}$ such that $\Phi_{i_{j}}^{-1}(t_{j-1})\in U$ and letting $t_{j}=\Phi_{i_{j}}^{-1}(t_{j-1})$. At the moment when $s\in B(t_{k},\varepsilon/M^{-k})$ we stop.
We note that once $k$ is such that $\varepsilon M^{-k}\geq 2r$ we have $s\in B(t_{k},\varepsilon/M^{-k})$, since both $t_{k},s\in U=B(m,r)$. This implies part (i).
To see parts (ii) and (iii). Let us denote $s_{k}=s$ and for $j=k-1,\ldots,0$ let $s_{j}=\Phi_{i_{j+1}}(s_{j+1}).$ We claim that for all $j=0,\ldots,k$ we have $|s_{j}-t_{j}|\leq 2r$ implying $s_{j}\in B(m,3r)$ as well as $|s_{j}-t_{j}|\leq\varepsilon/M^{-j}$. Together this implies parts (ii) and (iii). We prove the claim by descending induction. For $j=k$ the two claims are true by construction. Now assume $j<k$. Since $s_{j+1},t_{j+1}$ are both contained in $B(m,3r)$, we have by the fundamental theorem of calculus
$|s_{j}-t_{j}|=|\Phi_{i_{j+1}}(s_{j+1})-\Phi_{i_{j+1}}(t_{j+1})|\leq\sup_{z\in B(m,3r)}|\Phi_{i_{j+1}}^{\prime}(z)|\cdot|s_{j+1}-t_{j+1}|\leq M|s_{j+1}-t_{j+1}|.$
So by induction we have $|s_{j}-t_{j}|\leq\varepsilon M^{-j}$ and $|s_{j}-t_{j}|\leq M2r<2r$. This finishes the proof.
A similar proof also works in the special case when all parameters are real. ∎
@@
By Theorem 3.5 we have $p\in\mathcal{A}_{G_{0}}$ (resp. $p\in\mathcal{A}^{\mathbb{R}}_{G_{0}}$) and thus by Lemma 3.4 there exists $\alpha\in(0,1)$ such that $|g^{\prime}(1)|<1-\alpha$.
@@
Consider the open set $U_{1}=\{z\in\mathbb{C}\mid|g^{\prime}(z)|<1-\alpha/2\}$. Clearly, $1\in U_{1}$. Since $U_{1}$ is open there exists $r>0$ such that $B(1,3r)\subseteq U_{1}$. (We can easily determine such $r$ explicitly from the formula for $g^{\prime}(z)$.) Then for every $u\in B(1,3r)$ we have that $z\mapsto g(z)+u-1$ has the absolute value of its derivative bounded by $1-\alpha/2$ on $B(1,3r)$. Let $U=B(1,r)$. Since $\{y_{G}(p)\mid G\in\mathcal{H}_{G_{0}},R(G;p)\neq 0\}$ is dense in $\mathbb{C}$ the collection of open sets of the form $g(U)+y_{G}(p)-1$ with $G\in\mathcal{H}^{*}_{G_{0}}(p)$ covers the compact set $\overline{U}$. Therefore there exists a finite set of graphs $G_{i}\in\mathcal{H}^{*}_{G_{0}}(p)$, $i=1,\ldots,t$ such the associated sets cover $\overline{U}$.
Therefore, the corresponding set of contracting Möbius transformations $\{\Phi_{i}(z)=g(z)+y_{G_{i}}(p)-1\}_{i\in[t]}$ satisfies the hypothesis of Lemma 5.3. ∎
@@
Now we are ready to describe the desired algorithm. Let $\omega_{0}=y_{0}+(1+p)$. Our goal is to find an effective edge interaction $\omega_{1}=y_{G}(p)$ such that $|\omega_{1}-\omega_{0}|\leq\varepsilon$ in polynomial time of the input. For the algorithm we consider three different cases: (1) $\omega_{0}\in U$, (2) $\omega_{0}\in B(p,2)\setminus U$ and (3) $\omega_{0}\in\mathbb{C}\setminus(U\cup B(p,2))$.